The weights of adult giraffes follow a normal distribution with mean 2200 pounds and standard deviation 200 pounds. What is the probability that a randomly selected adult giraffe weighs more than 2350 pounds? a) 0.227 b) 0.273 c) 0.469 d) 0.518 e) 0.773
Answer #1

Solution :

Given ,

mean = = 2200

standard deviation = = 200

P(x >2350 ) = 1 - P(x<2350 )

= 1 - P[(x -) / < (2350-2200) / 200]

= 1 - P(z <0.75 )

Using z table

= 1 - 0.7734

= 0.2266

probability=0.227